# DARTH: Declarative Recall Through Early Termination for Approximate Nearest Neighbor Search

### v1

@@ Abstract.
Approximate Nearest Neighbor Search (ANNS) presents an inherent tradeoff between performance and recall (i.e., result quality). Each ANNS algorithm provides its own algorithm-dependent parameters to allow applications to influence the recall/performance tradeoff of their searches. This situation is doubly problematic. First, the application developers have to experiment with these algorithm-dependent parameters to fine-tune the parameters that produce the desired recall for each use case. This process usually takes a lot of effort. Even worse, the chosen parameters may produce good recall for some queries, but bad recall for hard queries. To solve these problems, we present DARTH, a method that uses target declarative recall. DARTH uses a novel method for providing target declarative recall on top of an ANNS index by employing an adaptive early termination strategy integrated into the search algorithm. Through a wide range of experiments, we demonstrate that DARTH effectively meets user-defined recall targets while achieving significant speedups, up to 14.6x (average: 6.8x; median: 5.7x) faster than the search without early termination for HNSW and up to 41.8x (average: 13.6x; median: 8.1x) for IVF. This paper appeared in ACM SIGMOD 2026.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Motivation. Approximate Nearest Neighbor Search (ANNS) for high-dimensional vector databases (Echihabi et al., 2020b; Wang et al., 2023a) is heavily used for semantic search in multiple application areas (Echihabi et al., 2020a), including web search engines (Chen et al., 2021; Chatzakis et al., 2021), multimedia databases (Ferhatosmanoglu et al., 2001; Sanca et al., 2023), recommendation systems (Das et al., 2007; Chen et al., 2022; Rajput et al., 2023), image retrieval (Wei et al., 2020; Ye and Xu, 2003), Large Language Models (LLM) (AI, 2023; OpenAI, 2024; Touvron et al., 2023) and Retrieval Augmented Generation (RAG) (Gao et al., 2023; Li et al., 2022; Lewis et al., 2020). ANNS has attracted massive industrial interest recently as new generations of embedding models have enabled powerful semantic search. In response, multiple SQL and NoSQL database vendors have recently announced ANN indices in support of ANNS (pgv, 2024; Cloud, 2024; Microsoft Azure, [n. d.]; MongoDB, Inc., [n. d.]; Elastic, [n. d.]; Oracle Corporation, [n. d.]; Amazon Web Services, [n. d.]) and, furthermore, multiple purpose-built vector databases featuring ANNS have been launched by startups (Pinecone, Inc., [n. d.]; Technologies, 2019; Wang et al., 2021b) and from cloud providers (Google Cloud, [n. d.]; Microsoft, [n. d.]).
@@
We evaluate the efficiency of DARTH through an extensive experimental evaluation using 5 popular datasets of varying sizes and dimensionalities. Our results demonstrate that the early termination recall of DARTH is accurate: DARTH is always able to meet the user-declared recall targets while offering significant speedups. Specifically, we show that our approach achieves up to 14.6x (average 6.8x, median 5.7x) speedup compared to the HNSW search without early termination. DARTH terminates the search very near the optimal point, performing on average only 5% more distance calculations than the optimal. We compare our approach to several other approaches for declarative recall, and we show that DARTH provides State-of-the-Art (SotA) search quality results, while delivering efficient search times. We show the superiority of DARTH for query workloads that include harder and Out-Of-Distribution (OOD) queries, demonstrating that DARTH is the method that achieves the best results. Lastly, we demonstrate that DARTH is efficient for IVF as well, always meeting the declared recall targets and achieving up to 41.8x (average 13.6x, median 8.1x) speedup compared to IVF search without early termination.
@@ 2.1. Preliminaries
ANNS Indices. ANNS tasks are efficiently addressed using specialized ANNS indices (Wang et al., 2023a; Azizi et al., 2025). These approaches construct an index structure over the vector collection $V$, enabling rapid query answering times. Such indices generally fall into four main categories: Tree-based (Yagoubi et al., 2017; Palpanas, 2020; Yagoubi et al., 2020; Echihabi et al., 2022; Chatzakis et al., 2023; Wang et al., 2023b; Peng et al., 2020, 2021; Wang et al., 2024b; Fatourou et al., 2023; Wang et al., 2025), LSH-based (Huang et al., 2015; Dasgupta et al., 2011), Quantization-based (Gao and Long, 2024; Matsui et al., 2018; Ge et al., 2013), Graph-based (Wang et al., 2021a, 2023a; Fu et al., 2017; Jayaram Subramanya et al., 2019; Gollapudi et al., 2023; Gou et al., 2024). In addition, several hybrid methods have emerged, such as ELPIS (Azizi et al., 2023) (Tree-Graph-based), DET-LSH (Wei et al., 2024) (Tree-LSH-based), ScaNN (Guo et al., 2020) and IVF-PQ (Jegou et al., 2010a; Douze et al., 2024) (Tree-Quantization-based), and others (Chen et al., 2021; Doshi et al., 2020; Wei et al., 2025). Graph-based indices, which are the primary focus of this work, create a graph over $V$ by representing vectors as nodes, with edges between them reflecting some measure of proximity between the nodes. There are numerous variations in graph-based methods, such as HNSW (Malkov and Yashunin, 2018), DiskANN (Jayaram Subramanya et al., 2019) and others (Dong et al., 2011; Fu et al., 2017; Gou et al., 2024). Still, the search process for a query remains largely consistent between all approaches since the main operation is to traverse the graph, collecting the nearest neighbors of a query.
@@
Hierarchical Navigable Small World (HNSW) graph. The HNSW graph (Malkov and Yashunin, 2018) is one of the most efficient and accurate SotA indices for ANNS (Wang et al., 2023a; Aumüller et al., 2020; Wang et al., 2021a). It organizes vectors into a multi-layered hierarchical structure, where each layer represents different levels of proximity. Vectors are inserted starting from the base (lowest) layer, with higher layers being created probabilistically. The key parameters that influence the performance of HNSW graph creation are $M$, $efConstruction$, and $efSearch$. The parameter $M$ defines the maximum number of neighbors a vector can have. A higher value of $M$ improves search quality by making the graph denser, but it also increases memory usage and search time. The parameter $efConstruction$ controls the number of candidates considered during graph construction, with larger values resulting in a more accurate graph at the cost of longer construction times. An overview of the query phase is illustrated in Figure 2(a). The search for a query starts from the top layer of the graph, from a predefined entry point. The search progresses greedily, progressively using the closest node of each layer as an entry point for the next layer, until the base layer of the graph (which contains all vectors of the dataset) is reached. Once the search reaches the base layer, it continues with a detailed traversal of candidate neighbors (shown in green) to retrieve the most similar vectors, putting the candidate vectors in a priority queue, and by putting the collected nearest neighbors in a result set, usually implemented as a heap. The amount of search effort in the base layer is influenced by the parameter $efSearch$, which determines the number of candidate neighbors to examine during query processing. A higher $efSearch$ leads to better recall but at the expense of longer search times. The HNSW search in the base layer terminates when no better candidates remain to be added to the priority queue—meaning all vectors in the priority queue are closer to the query than their unexplored neighbors—or when the entire base layer has been searched (a very rare occurrence). These termination points, occurring without early termination, are referred to as natural termination points, and the HNSW index that employs the search algorithm described above, terminating at the natural termination points is referred to as plain HNSW.
@@
Figure 2. (a): Example of locating the nearest neighbor of a query in an HNSW Graph. (b): Algorithms $A$ and $B$ achieve the same recall, yet, the algorithm $A$ results are of higher quality.
@@ 2.2. Related Work
Early Termination Approaches. To the best of our knowledge, DARTH is the only approach that directly and natively tackles the problem of declarative recall using early termination. Recently, early termination techniques for ANNS have been proposed. These methods aim to terminate the search for a query as soon as a specific algorithm-specific objective is met (e.g., all nearest neighbors are found), thus improving search time. The current SotA approaches are ProS (Gogolou et al., 2019; Echihabi et al., 2023) and Learned Adaptive Early termination (Li et al., 2020). Both approaches leverage the observation that, in nearest neighbor search (both exact and approximate), the k nearest neighbors of a query are typically found early in the search process, allowing a significant portion of the search to be skipped. ProS employs statistical and Machine Learning (ML) models to terminate the search early once all nearest neighbors are found, focusing on exact similarity search for Data Series using the iSAX (Camerra et al., 2010) index. It is a progressive approach, meaning that during the search for the neighbors of a query, the model is utilized multiple times to decide if all nearest neighbors are found, allowing for progressively better and more accurate predictions. In contrast, Learned Adaptive Early Termination uses an ML model to predict how many distance calculations are required for a query to retrieve all nearest neighbors that the index search algorithm would find, targeting the HNSW and IVF-PQ (Jegou et al., 2010a) indices. In this method, the model is called only once at a specific time during the search, indicating the total number of distance calculations that need to be performed.
@@ 3.1.1. Descriptive Input Features
$\bullet$ Index features: These features provide insight into the progression of the search process. They include the current step of the search conducted at the base layer of the HNSW at the time of observation ($nstep$), the number of distance calculations performed ($ndis$), and the number of updates to the nearest neighbor result set up to that point ($ninserts$).
@@
$\bullet$ Nearest Neighbor (NN) Stats features: These features provide descriptive summary statistics of the nearest neighbors found for the query up to a given point in the search. They include the average ($avg$), the variance ($var$), the median ($med$), and the 25th and 75th percentiles ($perc25$, $perc75$) of the nearest neighbor distances in the result set.
@@
Type Features Description Index $nstep$ Search step $ndis$ No. distance calculations $ninserts$ No. updates in the NN result set NN Distance $firstNN$ Distance of first NN found $closestNN$ Distance of current closest NN $furthestNN$ Distance of current furthest ($k$-th) NN NN Stats $avg$ Average of distances of the NN $var$ Variance of distances of NN $med$ Median of distances of NN $perc25$ 25th percentile of distances of NN $perc75$ 75th percentile of distances of NN
@@ 3.2.1. Adaptive Prediction Interval
where $pi$ is the new (updated) prediction interval, $mpi$ is the minimum prediction interval allowed, $ipi$ is the initial prediction interval (the recall predictor will be called for the first time after $ipi$ distance calculations), $R_{t}$ is the target recall and $R_{p}$ is the predicted recall as predicted from the model. This linear formula generates smaller prediction intervals when $R_{p}$ is close to $R_{t}$, and larger prediction intervals when $R_{p}$ is far from $R_{t}$.
@@ 3.2.2. Hyperparameter Importance
The introduction of two hyperparameters, $ipi$ (initial/max prediction interval) and $mpi$ (minimum prediction interval), is a crucial aspect of our approach. These hyperparameters control how frequently the predictor is called, with $pi\in[mpi,ipi]$. Setting appropriate values for these hyperparameters is essential: for instance, a very high value for $ipi$ may delay the initial predictor call, missing early opportunities for termination, while a very low value for $mpi$ could lead to an excessive number of predictor invocations, thereby introducing unnecessary overhead. The values for the hyperparameters can be selected either by classic grid-search tuning (or other sophisticated hyperparameter tuning approaches) or by a generic, heuristic-based selection method. For the generic heuristic-based method, to find a suitable value of $ipi$ for a specific recall target $R_{t}$, we calculate the average number of distance calculations needed to reach this target from the training queries, denoted as $dists_{R_{t}}$. This information is readily available during the generation of training data from our training queries, incurring no additional costs. We then set the values for our hyperparameters as $ipi=\frac{dists_{R_{t}}}{2}$ and $mpi=\frac{dists_{R_{t}}}{10}$ In addition, this method imposes an interesting baseline for comparison to our approach. In our experimental evaluation (Section 4), we analyze several aspects of hyperparameter selection, including the superiority of adaptive intervals compared to static intervals, as well as the comparison between the generic heuristic selection approach and the extensively tuned selection approach. Our evaluation shows that the heuristic parameters result in a very close performance to that achieved with the extensively tuned parameters. This means that DARTH requires no hyperparameter tuning, which is a significant improvement over the available competitors. Also, our experimental evaluation compares DARTH against a Baseline for early termination which early terminates every HNSW search after $dists_{R_{t}}$ distance calculations for a recall target $R_{t}$, showing that this approach is not sufficient to solve our research problem.
@@ 3.3.1. Integration in HNSW
10:  HNSW Graph $G$, Query Vector $q$, Initial Prediction Interval $ipi$, Minimum Prediction Interval $mpi$, Recall Predictor $model$, Number of neighbors to return $k$, Target Recall $R_{t}$, Search effort parameter $efSearch$ 21:  Traverse the upper layers of $G$ with beam search width 1 to reach the base layer (BL) 32:  $N_{BL}$leftarrow Entry node of the base layer (BL) 43:  $firstNN\leftarrow\text{Distance}(q,N_{BL})$ 54:  Initialize $resultSet$ as a heap of size $k$ 5:  Initialize counters: $ndis,nstep,inserts\leftarrow 0$ 6:  Initialize counter: $idis\leftarrow 0$ 67:  Set initial prediction interval: $pi\leftarrow ipi$ 8:  Initialize priority queue $candidateQueue$ of size $efSearch$ 79:  Add ($N_{BL}$,$firstNN$) to $candidateQueue$ 10:  while $candidateQueue$ is not empty do 811:     Extract node $c$ from $candidateQueue$ with the minimum distance 912:     $ndis\leftarrow ndis+1$, $idis\leftarrow idis+1$ 13:     Compute $cDis\leftarrow\text{Distance}(q,c)$ 14:     if $cDis<\text{GetMaxDistance}(resultSet)$ then 15:        Add $(c,cDis)$ to $resultSet$ 16:        $inserts\leftarrow inserts+1$ 1017:     end if 18:     for each unvisited neighbor node $n$ of $c$ do 1119:        Compute $nDis\leftarrow\text{Distance}(q,n)$ 20:        if $nDis<\text{GetMaxDistance}(resultSet)$ or $|candidateQueue|<efSearch$ then 21:           Add ($n$, $nDis$) to $candidateQueue$ 22:        end if 1223:     end for 24:     if $idis\bmod pi=0$ then 25:        Prepare input vector $input$ with features from Table 1 1326:        $R_{p}\leftarrow\text{model.predict}(input)$ 27:        if $R_{p}\geq R_{t}$ then 28:           return  $resultSet$ 1429:        end if 30:        Adjust prediction interval: $pi\leftarrow mpi+(ipi-mpi)\cdot(R_{t}-R_{p})$ 31:        Reset interval counter: $idis\leftarrow 0$ 1532:     end if 33:     $nstep\leftarrow nstep+1$ 1634:  end while 35:  return  $resultSet$

### v2

@@ Abstract.
Approximate Nearest Neighbor Search (ANNS) presents an inherent tradeoff between performance and recall (i.e., result quality). Each ANNS algorithm provides its own algorithm-dependent parameters to allow applications to influence the recall/performance tradeoff of their searches. This situation is doubly problematic. First, the application developers have to experiment with these algorithm-dependent parameters to fine-tune the parameters that produce the desired recall for each use case. This process usually takes a lot of effort. Even worse, the chosen parameters may produce good recall for some queries, but bad recall for hard queries. To solve these problems, we present DARTH, a method that uses target declarative recall. DARTH uses a novel method for providing target declarative recall on top of an ANNS index by employing an adaptive early termination strategy integrated into the search algorithm. Through a wide range of experiments, we demonstrate that DARTH effectively meets user-defined recall targets while achieving significant speedups, up to 14.6x (average: 6.8x; median: 5.7x) faster than the search without early termination for HNSW and up to 41.8x (average: 13.6x; median: 8.1x) for IVF.
This paper appeared in ACM SIGMOD 2026.
@@
copyright: ccjournal: PACMMODjournalyear: 2025journalvolume: 3journalnumber: 4 (SIGMOD)article: 242publicationmonth: 9doi: 10.1145/3749160ccs: Information systems Query optimization
@@ 1. Introduction
Motivation. Approximate Nearest Neighbor Search (ANNS) for high-dimensional vector databases (Echihabi et al., 2020b; Wang et al., 2023a) is heavily used for semantic search in multiple application areas (Echihabi et al., 2020a), including web search engines (Chen et al., 2021; Chatzakis et al., 2021), multimedia databases (Ferhatosmanoglu et al., 2001; Sanca et al., 2024), recommendation systems (Das et al., 2007; Chen et al., 2022; Rajput et al., 2023), image retrieval (Wei et al., 2020; Ye and Xu, 2003), Large Language Models (LLM) (AI, 2023; OpenAI, 2024; Touvron et al., 2023) and Retrieval Augmented Generation (RAG) (Gao et al., 2023; Li et al., 2022; Lewis et al., 2020). ANNS has attracted massive industrial interest recently as new generations of embedding models have enabled powerful semantic search. In response, multiple SQL and NoSQL database vendors have recently announced ANN indices in support of ANNS (pgv, 2024; Cloud, 2024; Microsoft Azure, [n. d.]; MongoDB, Inc., [n. d.]; Elastic, [n. d.]; Oracle Corporation, [n. d.]; Amazon Web Services, [n. d.]) and, furthermore, multiple purpose-built vector databases featuring ANNS have been launched by startups (Pinecone, Inc., [n. d.]; Technologies, 2019; Wang et al., 2021b) and from cloud providers (Google Cloud, [n. d.]; Microsoft, [n. d.]).
@@
We evaluate the efficiency of DARTH through an extensive experimental evaluation using 5 popular datasets of varying sizes and dimensionalities. Our results demonstrate that the early termination recall of DARTH is accurate: DARTH is always able to meet the user-declared recall targets while offering significant speedups. Specifically, we show that our approach achieves up to 14.6x (average 6.8x, median 5.7x) speedup compared to the HNSW search without early termination. DARTH terminates the search very near the optimal point, performing on average only 5% more distance calculations than the optimal. We compare our approach to several other approaches for declarative recall, and we show that DARTH provides State-of-the-Art (SotA) search quality results, while delivering efficient search times. We show the superiority of DARTH for query workloads that include harder and Out-Of-Distribution (OOD) queries, demonstrating that DARTH is the method that achieves the best results. Lastly, we demonstrate that DARTH is efficient for IVF as well, always meeting the declared recall targets and achieving up to 41.8x (average 13.6x, median 8.1x) speedup compared to IVF search without early termination. We make our code publicly available on GitHub (dar, 2025).
@@ 2.1. Preliminaries
ANNS Indices. ANNS tasks are efficiently addressed using specialized ANNS indices (Wang et al., 2023a; Azizi et al., 2025). These approaches construct an index structure over the vector collection $V$, enabling rapid query answering times. Such indices generally fall into four main categories: Tree-based (Yagoubi et al., 2017; Palpanas, 2020; Yagoubi et al., 2020; Echihabi et al., 2022; Chatzakis et al., 2023; Wang et al., 2023b; Peng et al., 2020, 2021; Wang et al., 2024b; Fatourou et al., 2023; Wang et al., 2025), LSH-based (Huang et al., 2015; Dasgupta et al., 2011), Quantization-based (Gao and Long, 2024; Matsui et al., 2018; Ge et al., 2013), Graph-based (Wang et al., 2021a, 2023a; Fu et al., 2017; Jayaram Subramanya et al., 2019; Gollapudi et al., 2023; Gou et al., 2024). In addition, several hybrid methods have emerged, such as ELPIS (Azizi et al., 2023) (Tree-Graph-based), DET-LSH (Wei et al., 2024) (Tree-LSH-based), ScaNN (Guo et al., 2020) and IVF-PQ (Jegou et al., 2010; Douze et al., 2024) (Tree-Quantization-based), and others (Chen et al., 2021; Doshi et al., 2020; Wei et al., 2025). Graph-based indices, which are the primary focus of this work, create a graph over $V$ by representing vectors as nodes, with edges between them reflecting some measure of proximity between the nodes. There are numerous variations in graph-based methods, such as HNSW (Malkov and Yashunin, 2018), DiskANN (Jayaram Subramanya et al., 2019) and others (Dong et al., 2011; Fu et al., 2017; Gou et al., 2024). Still, the search process for a query remains largely consistent between all approaches since the main operation is to traverse the graph, collecting the nearest neighbors of a query.
@@
Hierarchical Navigable Small World (HNSW) graph. The HNSW graph (Malkov and Yashunin, 2018) is one of the most efficient and accurate SotA indices for ANNS (Wang et al., 2023a; Aumüller et al., 2020; Wang et al., 2021a). It organizes vectors into a multi-layered hierarchical structure, where each layer represents different levels of proximity. Vectors are inserted starting from the base (lowest) layer, with higher layers being created probabilistically. The key parameters that influence the performance of HNSW graph creation are $M$, efConstructionefConstruction, and $efSearch$. The parameter $M$ defines the maximum number of neighbors a vector can have. A higher value of $M$ improves search quality by making the graph denser, but it also increases memory usage and search time. The parameter $efConstruction$ controls the number of candidates considered during graph construction, with larger values resulting in a more accurate graph at the cost of longer construction times. An overview of the query phase is illustrated in Figure 2(a). The search for a query starts from the top layer of the graph, from a predefined entry point. The search progresses greedily, progressively using the closest node of each layer as an entry point for the next layer, until the base layer of the graph (which contains all vectors of the dataset) is reached. Once the search reaches the base layer, it continues with a detailed traversal of candidate neighbors (shown in green) to retrieve the most similar vectors, putting the candidate vectors in a priority queue, and by putting the collected nearest neighbors in a result set, usually implemented as a heap. The amount of search effort in the base layer is influenced by the parameter $efSearch$, which determines the number of candidate neighbors to examine during query processing. A higher $efSearch$ leads to better recall but at the expense of longer search times. The HNSW search in the base layer terminates when no better candidates remain to be added to the priority queue—meaning all vectors in the priority queue are closer to the query than their unexplored neighbors—or when the entire base layer has been searched (a very rare occurrence). These termination points, occurring without early termination, are referred to as natural termination points, and the HNSW index that employs the search algorithm described above, terminating at the natural termination points is referred to as plain HNSW.
@@
Figure 2. (a): Example of locating the nearest neighbor of a query in an HNSW Graph. (b): Algorithms AA and $B$ achieve the same recall, yet, the algorithm AA results are of higher quality.
@@ 2.2. Related Work
Early Termination Approaches. To the best of our knowledge, DARTH is the only approach that directly and natively tackles the problem of declarative recall using early termination. Recently, early termination techniques for ANNS have been proposed. These methods aim to terminate the search for a query as soon as a specific algorithm-specific objective is met (e.g., all nearest neighbors are found), thus improving search time. The current SotA approaches are ProS (Gogolou et al., 2019; Echihabi et al., 2023) and Learned Adaptive Early termination (Li et al., 2020). Both approaches leverage the observation that, in nearest neighbor search (both exact and approximate), the k nearest neighbors of a query are typically found early in the search process, allowing a significant portion of the search to be skipped. ProS employs statistical and Machine Learning (ML) models to terminate the search early once all nearest neighbors are found, focusing on exact similarity search for Data Series using the iSAX (Camerra et al., 2010) index. It is a progressive approach, meaning that during the search for the neighbors of a query, the model is utilized multiple times to decide if all nearest neighbors are found, allowing for progressively better and more accurate predictions. In contrast, Learned Adaptive Early Termination uses an ML model to predict how many distance calculations are required for a query to retrieve all nearest neighbors that the index search algorithm would find, targeting the HNSW and IVF-PQ (Jegou et al., 2010) indices. In this method, the model is called only once at a specific time during the search, indicating the total number of distance calculations that need to be performed.
@@ 3.1.1. Descriptive Input Features
$\bullet$ Index features: These features provide insight into the progression of the search process. They include the current step of the search conducted at the base layer of the HNSW at the time of observation (nstepnstep), the number of distance calculations performed (ndisndis), and the number of updates to the nearest neighbor result set up to that point (ninsertsninserts).
@@
$\bullet$ Nearest Neighbor (NN) Stats features: These features provide descriptive summary statistics of the nearest neighbors found for the query up to a given point in the search. They include the average (avgavg), the variance (varvar), the median (medmed), and the 25th and 75th percentiles ($perc25$, $perc75$) of the nearest neighbor distances in the result set.
@@
Type Features Description Index nstepnstep Search step ndisndis No. distance calculations ninsertsninserts No. updates in the NN result set NN Distance $firstNN$ Distance of first NN found $closestNN$ Distance of current closest NN $furthestNN$ Distance of current furthest ($k$-th) NN NN Stats avgavg Average of distances of the NN varvar Variance of distances of NN medmed Median of distances of NN $perc25$ 25th percentile of distances of NN $perc75$ 75th percentile of distances of NN
@@ 3.2.1. Adaptive Prediction Interval
where $pi$ is the new (updated) prediction interval, mpimpi is the minimum prediction interval allowed, ipiipi is the initial prediction interval (the recall predictor will be called for the first time after ipiipi distance calculations), $R_{t}$ is the target recall and $R_{p}$ is the predicted recall as predicted from the model. This linear formula generates smaller prediction intervals when $R_{p}$ is close to $R_{t}$, and larger prediction intervals when $R_{p}$ is far from $R_{t}$.
@@ 3.2.2. Hyperparameter Importance
The introduction of two hyperparameters, ipiipi (initial/max prediction interval) and mpimpi (minimum prediction interval), is a crucial aspect of our approach. These hyperparameters control how frequently the predictor is called, with $pi\in[mpi,ipi]$. Setting appropriate values for these hyperparameters is essential: for instance, a very high value for ipiipi may delay the initial predictor call, missing early opportunities for termination, while a very low value for mpimpi could lead to an excessive number of predictor invocations, thereby introducing unnecessary overhead. The values for the hyperparameters can be selected either by classic grid-search tuning (or other sophisticated hyperparameter tuning approaches) or by a generic, heuristic-based selection method. For the generic heuristic-based method, to find a suitable value of ipiipi for a specific recall target $R_{t}$, we calculate the average number of distance calculations needed to reach this target from the training queries, denoted as $dists_{R_{t}}$. This information is readily available during the generation of training data from our training queries, incurring no additional costs. We then set the values for our hyperparameters as $ipi=\frac{dists_{R_{t}}}{2}$ and $mpi=\frac{dists_{R_{t}}}{10}$ In addition, this method imposes an interesting baseline for comparison to our approach. In our experimental evaluation (Section 4), we analyze several aspects of hyperparameter selection, including the superiority of adaptive intervals compared to static intervals, as well as the comparison between the generic heuristic selection approach and the extensively tuned selection approach. Our evaluation shows that the heuristic parameters result in a very close performance to that achieved with the extensively tuned parameters. This means that DARTH requires no hyperparameter tuning, which is a significant improvement over the available competitors. Also, our experimental evaluation compares DARTH against a Baseline for early termination which early terminates every HNSW search after $dists_{R_{t}}$ distance calculations for a recall target $R_{t}$, showing that this approach is not sufficient to solve our research problem.
@@ 3.3.1. Integration in HNSW
10:  HNSW Graph $G$, Query Vector $q$, Initial Prediction Interval ipiipi, Minimum Prediction Interval mpimpi, Recall Predictor modelmodel, Number of neighbors to return $k$, Target Recall $R_{t}$, Search effort parameter efSearchefSearch 21:  Traverse the upper layers of $G$ with beam search width 1 to reach the base layer (BL) 32:  $N_{BL}$ $\leftarrow$ Entry node of the base layer (BL) 43:  $firstNN\leftarrow\text{Distance}(q,N_{BL})$ 54:  Initialize resultSetresultSet as a heap of size $k$ 5:  Initialize counters: $ndis,nstep,inserts\leftarrow 0$ 6:  Initialize counter: $idis\leftarrow 0$ 67:  Set initial prediction interval: $pi\leftarrow ipi$ 8:  Initialize priority queue candidateQueuecandidateQueue of size efSearchefSearch 79:  Add ($N_{BL}$,firstNNfirstNN) to candidateQueuecandidateQueue 10:  while candidateQueuecandidateQueue is not empty do 811:   Extract node $c$ from candidateQueuecandidateQueue with the minimum distance 912:   $ndis\leftarrow ndis+1$, $idis\leftarrow idis+1$ 13:   Compute $cDis\leftarrow\text{Distance}(q,c)$ 14:   if $cDis<\text{GetMaxDistance}(resultSet)$ then 15:    Add $(c,cDis)$ to resultSetresultSet 16:    $inserts\leftarrow inserts+1$ 1017:   end if 18:   for each unvisited neighbor node $n$ of $c$ do 1119:    Compute $nDis\leftarrow\text{Distance}(q,n)$ 20:    if $nDis<\text{GetMaxDistance}(resultSet)$ or $|candidateQueue|<efSearch$ then 21:     Add ($n$, nDisnDis) to candidateQueuecandidateQueue 22:    end if 1223:   end for 24:   if $idis\bmod pi=0$ then 25:    Prepare input vector inputinput with features from Table 1 1326:    $R_{p}\leftarrow\text{model.predict}(input)$ 27:    if $R_{p}\geq R_{t}$ then 28:     return  resultSetresultSet 1429:    end if 30:    Adjust prediction interval: $pi\leftarrow mpi+(ipi-mpi)\cdot(R_{t}-R_{p})$ 31:    Reset interval counter: $idis\leftarrow 0$ 1532:   end if 33:   $nstep\leftarrow nstep+1$ 1634:  end while 35:  return  resultSetresultSet